# On the sample complexity of estimation in logistic regression

### v1

@@ Abstract
The logistic regression model is one of the most popular data generation model in noisy binary classification problems. In this work, we study the sample complexity of estimating the parameters of the logistic regression model up to a given $\ell_{2}$ error, in terms of the dimension and the inverse temperature, with standard normal covariates. The inverse temperature controls the signal-to-noise ratio of the data generation process. While both generalization bounds and asymptotic performance of the maximum-likelihood estimator for logistic regression are well-studied, the non-asymptotic sample complexity that shows the dependence on error and the inverse temperature for parameter estimation is absent from previous analyses. We show that the sample complexity curve has two change-points (or critical points) in terms of the inverse temperature, clearly separating the low, moderate, and high temperature regimes.
@@ 1 Introduction
The sample complexity $n^{*}(d,\beta,\epsilon)$ is the smallest sample size $n$ such that the above task is achievable by some estimator. We show that, up to logarithmic factors in $1/\epsilon$ in the expressions below, the sample complexity satisfies
$n^{*}(d,\beta,\epsilon)\asymp\begin{cases}\displaystyle\frac{d}{\beta^{2}\epsilon^{2}}&\text{if $\beta\lesssim 1$ (high temperatures)};\\ \displaystyle\frac{d}{\beta\epsilon^{2}}&\text{if $1\lesssim\beta\lesssim 1/\epsilon$ (moderate temperatures)};\\ \displaystyle\frac{d}{\epsilon}&\text{if $\beta\gtrsim 1/\epsilon$ (low temperatures)}.\end{cases}$
@@ 1.1 Motivation
improving on an earlier analysis of Plan and Vershynin (2012) that had a worse dependence on $\epsilon$; note that here we do not assume that $\theta^{\star}$ is sparse. The optimality of this estimator in the high-temperature regime ($\beta\lesssim 1$) is readily established as a standard application of Fano’s inequality (see, e.g., Chen et al., 2016, Appendix C.1). However, it was unclear whether this sample complexity is optimal in other regimes. In particular, in the zero-temperature regime (i.e., $\beta=\infty$), the $\mathbf{y}_{i}$ are deterministically given by the sign of $\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$, so the problem becomes equivalent to that of PAC learning homogeneous linear classifiers under spherically symmetric distributions on $\mathbf{x}$. For that problem, the sample complexity is $\Theta(d/\epsilon)$, as established by Long (1995, 2003); note that the dependence on $\epsilon$ is considerably reduced. [^1] Servedio (1999) studied the “Average” algorithm in this context, allowing for the possibility that each observed label is independently flipped with some fixed probability $\eta\in[0,1/2\rparen$ ), and obtains a sample complexity upper bound of $O(d/((1-2\eta)\epsilon)^{2})$. Jacques et al. (2013) also gives a statement of the lower bound in the context of one-bit compressive sensing.
@@
Our upper and lower bounds resolve the dependence of the sample complexity on the inverse temperature (up to logarithmic factors in $1/\epsilon$), particularly in the regime where $1\lesssim\beta<\infty$.
@@ 1.2 Techniques
Our lower bounds for the high and moderate temperature regimes (Theorem 1) are proved using Fano’s inequality; for convenience, we use a Bayesian version due to Zhang (2006), although the “Generalized Fano” approach of Han and Verdú (1994) would also work. The bound in the high temperature regime is a “textbook” application that uses a uniform quadratic bound on the Kullback–Leibler (KL) divergence between the distributions determined by nearby parameter vectors. However, the moderate temperature regime uses a more refined analysis that does not appear to be standard. To facilitate the required computation, we use the Bregman divergence form of the KL divergence between Bernoulli distributions.
@@
To establish upper bounds on the sample complexity, it is natural to consider the maximum likelihood estimator (MLE), which is equivalent to finding $\theta\in S^{d-1}$ that minimizes the empirical risk with respect to the logistic loss: $(1/n)\sum_{i=1}^{n}[g(\beta\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta)-\mathds{1}\{\mathbf{y}_{i}=1\}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta]$. The convexity of the logistic loss makes the MLE computationally tractable (perhaps after extending the parameter space to the ball). The risk of a given $\theta$ is the expected value of this empirical risk, and the excess risk relative to that of $\theta^{\star}$ is the KL divergence between the distributions determined by $\theta$ and $\theta^{\star}$. This KL divergence can, in turn, be related to the parameter error $\lVert\theta-\theta^{\star}\rVert$ using our analysis from the lower bound. However, bounding the excess risk sharply enough appears to be challenging. A standard approach in statistical learning theory is to use techniques like Rademacher averages from the theory of empirical processes to relate excess risks to the excess empirical risks. Unfortunately, using such tools designed specifically for smooth loss functions like the logistic loss (e.g., Srebro et al., 2010) leads to a sample complexity with suboptimal dependences on $\epsilon$ and $\beta$.
A more promising approach is to directly analyze the MLE by computing tight Taylor expansions of the estimation error in a way that keeps track of the dimension dependence (e.g., He and Shao, 2000; Portnoy, 1988), and, ideally, the inverse temperature. The most relevant work along these lines is that of Ostrovskii and Bach (2021) which, building on the fixed-design analysis of Bach (2010), leverages self-concordance properties of the logistic loss to directly bound the excess risk. [^2] The analysis of Bach (2010) and Ostrovskii and Bach (2021) can also apply to the setting where the logistic regression model is misspecified, but we only consider their analysis in the case where the model is well-specified. Although their analysis of the MLE matches the performance guarantees in the leading order terms as predicted by classical asymptotic analysis, it appears to be unable to avoid suboptimal dependence on the $\beta$ in the lower order terms.
We ultimately resort to establishing the upper bounds (Theorem 3), in all regimes, using empirical risk minimization with the classification loss, together with a precise analysis of the relationship between excess error rates and parameter distance. While this does not suggest any computationally tractable estimator (except, as far as we know, in the $\beta=\infty$ regime), the analysis is tight enough to match the lower bounds up to logarithmic factors in $1/\epsilon$.
Bounds on various Gaussian integrals used in our proofs are collected in Appendix A.
@@ 1.3 Other related works
Several prior works analyze “improper learning” algorithms for (possibly misspecified) logistic regression (e.g., Kakade and Ng, 2004; Zhang, 2006; Hazan et al., 2014; Foster et al., 2018; Mourtada and Gaïffas, 2022) that do not necessarily produce an estimate of $\theta^{\star}$, which may not be sensible anyway if the model is misspecified. Instead, the goal of these algorithms is to achieve low prediction error guarantees.
Probit model misspecified as logistic.
As we were preparing this manuscript, we became aware of the recent work of Kuchelmeister and van de Geer (2023), which studies the estimation of the parameter vector $\theta^{\star}\in S^{d-1}$ in a probit regression model
$\mathbf{y}=\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}+\sigma\mathbf{z}),\quad(\mathbf{x},\mathbf{z})\sim\operatorname{N}(0,I_{d})\otimes\operatorname{N}(0,1),$
by maximizing the likelihood under a (misspecified) logistic regression model. Their analysis reveals two distinct regimes for the rate of convergence, depending on how the noise level $\sigma>0$ compares to the sample size and dimension. (They also consider estimation of the norm of the expected likelihood maximizer, which is related to the signal-to-noise ratio.) Although their data model and the high-level findings are related to those of our work, we note that the aim of our work is to understand (upper and lower bounds on) sample complexity in a well-specified setup, whereas Kuchelmeister and van de Geer study a specific estimator based on a misspecified model.
@@ 1.4 Notations
For notational convenience, we assume that a Bernoulli distribution $\operatorname{Bern}(\mu)$ has support on $\{-1,1\}$, with the mean parameter $\mu$ being the probability of $1$. We will, on occasion, associate each $\theta\in S^{d-1}$ with a homogeneous linear classifier $h_{\theta}\colon\mathbb{R}^{d}\to\{-1,1\}$, given by $h_{\theta}(x)=1$ if $x^{\scriptscriptstyle{\mathsf{T}}}\theta>0$ and $h_{\theta}(x)=-1$ otherwise. For any $\theta,\theta^{\prime}\in S^{d-1}$, let $\operatorname{err}_{\theta}(\theta^{\prime})=\Pr(h_{\theta^{\prime}}(\mathbf{x})\neq\mathbf{y})=\Pr(\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\prime}\leq 0)$ be the error rate of $h_{\theta^{\prime}}$ when the distribution of $(\mathbf{x},\mathbf{y})$ is specified by parameter $\theta$.
@@ 2.1 Moderate and high temperatures
The proof of Theorem 1 uses the following information theoretic lower method due to Zhang (2006).
@@
Let $\Pi$ be a probabiltiy measure on a parameter space $\Theta$ indexing a family of probability measures $(P_{\theta})_{\theta\in\Theta}$ on a data space $\mathcal{Z}$, and let $L\colon\Theta\times\Theta\to\mathbb{R}$ be a loss function. Let $\boldsymbol{\theta}\sim\Pi$ and $\mathbf{Z}\mid\boldsymbol{\theta}\sim P_{\boldsymbol{\theta}}$. For any (possibly randomized) estimator ${\hat{\theta}}\colon\mathcal{Z}\to\Theta$,
@@
where we have used the triangle inequality $\lVert\theta-\theta^{\prime}\rVert\leq\lVert\theta-u\rVert+\lVert\theta^{\prime}-u\rVert\leq 2\delta$. By Lemma 10,
@@
$2\kappa+\ln 4\leq 2n\delta^{2}\beta\min\{\beta,2\sqrt{2/\pi}\}+\ln 4<(d-1)\ln 2$
as required, where the second inequality uses the assumption on $n$. ∎
@@ 2.2 Low temperatures
$\operatorname{err}_{\theta^{\star}}(\theta^{\star})\leq\min\left\{\frac{1}{2},\ \frac{1}{2}-\frac{\beta}{4}\sqrt{\frac{2}{\pi}}\left\lparen 1-\frac{\beta^{2}}{6}\right\rparen,\ \frac{1}{\beta}\sqrt{\frac{2}{\pi}}\right\}.$
@@
for a normal random variable $\mathbf{z}$, the claim follows immediately from Lemma 9. ∎
@@
$\operatorname{\mathbb{E}}\left[\sum_{\theta\in U}G_{\theta}\right]\geq\sum_{\theta\in U}\operatorname{\mathbb{E}}[G_{\theta}]\geq\sum_{\theta\in U}\left\lparen 1-\frac{1}{2}-\frac{1}{4}\right\rparen=\frac{\lvert U\rvert}{4}\geq\frac{1}{4}\left\lparen\frac{1}{2\epsilon}\right\rparen^{d-1}.$
@@ 3 Upper bounds on the sample complexity
We complement the sample complexity lower bounds from Theorems 1 and 2 with the following upper bound.

### v3

# On the sample complexity of parameter estimation in logistic regression with normal design

@@ Abstract
The logistic regression model is one of the most popular data generation model in noisy binary classification problems. In this work, we study the sample complexity of estimating the parameters of the logistic regression model up to a given $\ell_{2}$ error, in terms of the dimension and the inverse temperature, with standard normal covariates. The inverse temperature controls the signal-to-noise ratio of the data generation process. While both generalization bounds and asymptotic performance of the maximum-likelihood estimator for logistic regression are well-studied, the non-asymptotic sample complexity that shows the dependence on error and the inverse temperature for parameter estimation is absent from previous analyses. We show that the sample complexity curve has two change-points in terms of the inverse temperature, clearly separating the low, moderate, and high temperature regimes.
@@ 1 Introduction
The sample complexity $n^{\star}(d,\beta,\epsilon)$ is the smallest sample size $n$ such that the above task is achievable by some estimator. Intuitively, the estimation error for the maximum likelihood estimator here is asymptotically normal; leading one to believe that the sample complexity must scale as $d/\epsilon^{2}$. However, as $\beta$ increases to +$\infty$+ ∞, the estimation problem gradually becomes a noiseless halfspace learning problem. The sample complexity for the latter problem is known to be $d/\epsilon$. This means that the inverse temperature $\beta$ should play a crucial role in the sample complexity; going from small $\beta$ to large $\beta$, the sample complexity curve must have one or more change-points. In this paper, we coarsely determine this dependence on $\beta$. In particular, we show that, up to logarithmic factors in $d$ and $1/\epsilon$ in the expressions below, the sample complexity satisfies
$n^{\star}(d,\beta,\epsilon)\asymp\begin{cases}\displaystyle\frac{d}{\beta^{2}\epsilon^{2}}&\text{if $\beta\lesssim 1$ (high temperatures)};\vspace{1mm}\\ \displaystyle\frac{d}{\beta\epsilon^{2}}&\text{if $1\lesssim\beta\lesssim 1/\epsilon$ (moderate temperatures)};\vspace{1mm}\\ \displaystyle\frac{d}{\epsilon}&\text{if $\beta\gtrsim 1/\epsilon$ (low temperatures)}.\end{cases}$
@@ 1.1 Motivation
improving on an earlier analysis of Plan and Vershynin (2012) that had a worse dependence on $\epsilon$; note that here we do not assume that $\theta^{\star}$ is sparse. The optimality of this estimator in the high-temperature regime ($\beta\lesssim 1$) is readily established as a standard application of Fano’s inequality (see, e.g., Chen et al., 2016, Appendix C.1). However, it was unclear whether this sample complexity is optimal in other regimes. In particular, in the zero-temperature regime (i.e., $\beta=+\infty$), the $\mathbf{y}_{i}$ are determined by the sign of $\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$, so the problem becomes equivalent to that of PAC learning homogeneous linear classifiers under spherically symmetric distributions on $\mathbf{x}$. For that problem, the sample complexity is $\Theta(d/\epsilon)$, as established by Long (1995, 2003); note that the dependence on $\epsilon$ is considerably reduced. [^1] Servedio (1999) studied the “Average” algorithm in this context, allowing for the possibility that each observed label is independently flipped with some fixed probability $\eta\in[0,1/2\rparen$ ), and obtains a sample complexity upper bound of $O(d/((1-2\eta)\epsilon)^{2})$. Jacques et al. (2013) also gives a statement of the lower bound in the context of one-bit compressive sensing.
@@
Our upper and lower bounds resolve the dependence of the sample complexity on the inverse temperature (up to logarithmic factors in $1/\epsilon$ and $d$), particularly in the regime where $1\lesssim\beta<+\infty$.
@@ 1.2 Techniques
Lower bounds.
Our lower bounds for the high and moderate temperature regimes (Theorem 1) are proved using Fano’s inequality; for convenience, we use a Bayesian version due to Zhang (2006), although the “Generalized Fano” approach of Han and Verdú (1994) would also work. The bound in the high temperature regime is a “textbook” application that uses a uniform quadratic bound on the Kullback–Leibler (KL) divergence between the distributions determined by nearby parameter vectors. However, the moderate temperature regime requires a more refined analysis that does not appear to be standard. To facilitate the required computation, we use the Bregman divergence form of the KL divergence between Bernoulli distributions.
@@
To establish upper bounds on the sample complexity, it is natural to consider the maximum likelihood estimator (MLE), which is equivalent to finding $\theta\in S^{d-1}$ that minimizes the empirical risk with respect to the logistic loss: $(1/n)\sum_{i=1}^{n}g(-\beta\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta)$. The convexity of the logistic loss potentially makes the MLE computationally tractable (perhaps after extending the parameter space to the ball). The risk of a given $\theta$ is the expected value of this empirical risk, and the excess risk relative to that of $\theta^{\star}$ is the KL divergence between the distributions determined by $\theta$ and $\theta^{\star}$. This KL divergence can, in turn, be related to the parameter error $\lVert\theta-\theta^{\star}\rVert$ using our analysis from the lower bound. However, bounding the excess risk sharply enough appears to be challenging. A standard approach in statistical learning theory is to use techniques like Rademacher averages from the theory of empirical processes to relate excess risks to the excess empirical risks. Unfortunately, using such tools designed for smooth loss functions like the logistic loss (e.g., Srebro et al., 2010) leads to a sample complexity with suboptimal dependences on $\epsilon$ and $\beta$. Although the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ satisfies the Tsybakov-Mammen margin condition $\Pr[\lvert g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})-1/2\rvert\leq t]\leq C_{0}t^{\alpha/(1-\alpha)}$ with $C_{0}=O(\beta)$ and $\alpha=1/2$ (Mammen and Tsybakov, 1999), the improved $n^{-2/3}$ rate for the excess classification risk comes with a prefactor that grows with $\beta^{1/3}$.
A different approach is to directly analyze the MLE by computing tight Taylor expansions of the estimation error, and using properties such as self-concordance of the logistic function, in a way that keeps track of the dimension dependence (e.g., He and Shao, 2000; Portnoy, 1988; Bach, 2010; Ostrovskii and Bach, 2021), and, ideally, the inverse temperature. These analyses of the MLE match the performance guarantees in the leading order terms as predicted by classical asymptotic analysis. However, they essentially treat $\beta$ as a constant (resulting in suboptimal dependence on $\beta$ in lower order terms), which we cannot afford to do as the moderate and low temperature regimes are defined by the comparison of $\beta$ to $1/\epsilon$. Redoing the analysis entirely in our setting boils down to showing $(\theta-\theta^{\star})^{\scriptscriptstyle{\mathsf{T}}}\nabla L(\theta)>0$ for all $\theta$ that are $\epsilon$ away from the true parameter vector $\theta^{\star}$, with $L$ being the negative log-likelihood function. At this point, we need a tight approximation of $\nabla L(\theta)$ around $\nabla L(\theta^{\star})$, which one may hope to obtain using the self-concordance property. However, in our setting, the inner product depends on $\theta$, and a uniform bound does not give the optimal scaling.
A very recent work of Kuchelmeister and van de Geer (2023) studies the estimation of the parameter vector in a probit regression model (under a normal design) by maximizing the likelihood under a (misspecified) logistic regression model. They directly use techniques from empirical process theory specialized to the normal design to provide upper bounds on the sample complexity needed to estimate $\theta^{\star}$ up to a given $\ell_{2}$ error (under certain assumptions on the signal-to-noise ratio in the probit model; they are also concerned with estimation of the signal-to-noise ratio itself, which is beyond the scope of our work). We leverage essential parts of their analysis to establish our sample complexity upper bound in the moderate and low temperature regimes of our problem, which requires new moment bounds in the logistic regression observation model. We also give sample complexity bounds based on empirical risk minimization (for zero-one loss) in the low temperature regime, mostly using standard techniques. As mentioned above, an optimal estimator for the high temperature regime was already known (Plan et al., 2017).
Some of estimators we study are applicable under other designs, but the analyses in this work make heavy use of symmetry properties of the normal distribution. Bounds on various Gaussian integrals essential in our proofs are collected in Appendix A.
@@ 1.3 Other related works
Several prior works analyze “improper learning” algorithms for (possibly misspecified) logistic regression (e.g., Kakade and Ng, 2004; Zhang, 2006; Hazan et al., 2014; Foster et al., 2018; Mourtada and Gaïffas, 2022) that do not necessarily produce an estimate of $\theta^{\star}$, which may not be sensible anyway if the model is misspecified. Instead, the goal of these algorithms is to achieve low prediction error guarantees. The lower bound of Hazan et al. (2014) exploits misspecification to show that larger parameter norm is more detrimental to proper online learning than it is to improper online learning. In our well-specified setting for parameter estimation, the parameter norm has a very different effect.
High-dimensional proportional asymptotic analysis.
Another line of work considers the performance of MLE and regularized variants in the proportional asymptotic regime, where both $d$ and $n$ increase to infinity with $d/n$ tending to a constant $\delta>0$. In this setup, Sur and Candès (2019) are able to precisely characterize the region in the plane of $\delta$ and $\beta$ where the MLE exists. Salehi et al. (2019) consider the various regularized variants of MLE and characterize their asymptotic performance. Neither work directly reveals the dependence of the sample complexity on $\beta$.
@@ 1.4 Notations
For notational convenience, we assume that a Bernoulli distribution $\operatorname{Bern}(p)$ has support on $\{-1,1\}$, with the “mean parameter” $p$ being the probability of $1$. We occasionally associate each $\theta\in S^{d-1}$ with a homogeneous linear classifier $h_{\theta}\colon\mathbb{R}^{d}\to\{-1,1\}$, given by $h_{\theta}(x)=1$ if $x^{\scriptscriptstyle{\mathsf{T}}}\theta>0$ and $h_{\theta}(x)=-1$ otherwise. For any $\theta,\theta^{\prime}\in S^{d-1}$, let $\operatorname{err}_{\theta}(\theta^{\prime})=\Pr(h_{\theta^{\prime}}(\mathbf{x})\neq\mathbf{y})=\Pr(\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\prime}\leq 0)$ be the error rate of $h_{\theta^{\prime}}$ when the distribution of $(\mathbf{x},\mathbf{y})$ is specified by parameter $\theta$.
@@ 2.1 Moderate and high temperatures
The proof of Theorem 1 uses the following information theoretic lower bound method due to Zhang (2006).
@@
Let $\Pi$ be a probability measure on a parameter space $\Theta$ indexing a family of probability measures $(P_{\theta})_{\theta\in\Theta}$ on a data space $\mathcal{Z}$, and let $L\colon\Theta\times\Theta\to\mathbb{R}$ be a loss function. Let $\boldsymbol{\theta}\sim\Pi$ and $\mathbf{Z}\mid\boldsymbol{\theta}\sim P_{\boldsymbol{\theta}}$. For any (possibly randomized) estimator ${\hat{\theta}}\colon\mathcal{Z}\to\Theta$,
@@
where we have used the triangle inequality $\lVert\theta-\theta^{\prime}\rVert\leq\lVert\theta-u\rVert+\lVert\theta^{\prime}-u\rVert\leq 2\delta$. By Lemma 15 (see Appendix A),
@@
$2\kappa+\ln 4\leq 2n\delta^{2}\beta\min\{\beta,2\sqrt{2/\pi}\}+\ln 4<(d-1)\ln 2.\qed$
@@ 2.2 Low temperatures
$\operatorname{err}_{\theta^{\star}}(\theta^{\star})\leq\frac{1}{\beta}\sqrt{\frac{2}{\pi}}.$
@@
for a normal random variable $\mathbf{z}$, the claim now follows from Lemma 13 (see Appendix A). ∎
@@
$\operatorname{\mathbb{E}}\left[\sum_{\theta\in U}G_{\theta}\right]=\sum_{\theta\in U}\operatorname{\mathbb{E}}[G_{\theta}]\geq\sum_{\theta\in U}\left\lparen 1-\frac{1}{2}-\frac{1}{4}\right\rparen=\frac{\lvert U\rvert}{4}\geq\frac{1}{4}\left\lparen\frac{1}{2\epsilon}\right\rparen^{d-1}.$
@@ 3 Upper bounds on the sample complexity
In this section, we give upper bounds on the sample complexity based on three different estimators for the three different regimes of $\beta$. Throughout this section, we fix a “ground truth” parameter $\theta^{\star}\in S^{d-1}$ determining the distribution of $(\mathbf{x},\mathbf{y})$.
3.1 High temperatures
The analysis of Plan et al. (2017, Corollary 3.5) implies, for any $\epsilon\in\lparen 0,1\rparen$ ), the “linear estimator” ${\hat{\theta}_{\operatorname{linear}}}$ of Plan and Vershynin (2012) (or the “Average” algorithm of Servedio (1999))
${\hat{\theta}_{\operatorname{linear}}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})=\operatorname*{\arg\max}_{\theta\in S^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta$
satisfies $\lVert{\hat{\theta}_{\operatorname{linear}}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})-\theta^{\star}\rVert\leq\epsilon$ in expectation provided that
$n\geq C\max\left\{\frac{1}{\beta^{2}},1\right\}\frac{d}{\epsilon^{2}},$ (1)
where $C>0$ is a universal positive constant. This matches the lower bound from Theorem 1 (up to constants) when $\beta\lesssim 1$.
3.2 Moderate and low temperatures
When $\beta\gtrsim 1$, the sample size requirement of the linear estimator in Equation 1 has a suboptimal dependence on $\beta$. In particular, the sample size requirement does not become smaller as $\beta$ becomes larger.
In this section, we analyze a different estimator, the empirical ReLU risk minimizer ${\hat{\theta}_{\operatorname{relu}}}$:
${\hat{\theta}_{\operatorname{relu}}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})\in\operatorname*{\arg\min}_{\theta\in S^{d-1}}\frac{1}{n}\sum_{i=1}^{n}[-\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta]_{+},$ (2)
where $[x]_{+}=\max\{0,x\}$ (the rectified linear function used in ReLUs). Notice that if $[\cdot]_{+}$ was omitted in each summand of Equation 2, then the estimator ${\hat{\theta}_{\operatorname{relu}}}$ would be the same as ${\hat{\theta}_{\operatorname{linear}}}$. This estimator is also related to the Perceptron algorithm (Rosenblatt, 1958), since the latter can be viewed as a stochastic subgradient method for minimizing the empirical ReLU risk. However, we do not know if such subgradient methods minimize the objective over the nonconvex domain $S^{d-1}$.